A six-passenger plane cruises at 180 mph in calm air. If the plane flies 7 miles with the wind in the same amount of time as it flies 5 miles against the wind, then what is the speed of the wind?
Answer by checkley71(8403)

180/(x+5)=5
180/(x-5)=7
180/(x+5)+2=180/(x-5)
(180+2X+10)/(x+5)=180/(X-5)
(190+2X)/(X+5)=180/(X-5)
180(X+5)=(190+2X)(X-5)
180X+900=190X+2X^2-950-10X
2X^2+190X-10X-180X-950-900=0
2X^2-1850=0
X^2-925=0
(X-30.41)(X+30.41)=0
X-30.41=0
X=30.41 MPH FOR THE WIND.
PROOF
180/(30.42+5)=5
180/35.41=5
5=5
180/(30.41-5)=7
180/25.41=7
7=7